# Global well-posedness of 2D Navier-Stokes with Dirichlet boundary fractional noise

### v1

@@ Acknowledgements.
The authors thank Robert Denk and Tim Seitz for several useful discussions.
@@ 1. Introduction
In many situations occurring in applied sciences, noise can affect the evolution of a system only through the boundary of a region where the system evolves. Such phenomena can be modeled via partial differential equations with boundary noise, as introduced by Da Prato and Zabczyck in the seminal paper [20]. Such a description presents several issues from a mathematical viewpoint. Indeed, nowadays it is well-known that in the one dimensional case, the solution of the heat equation with white noise Dirichlet or Neumann boundary conditions has low (space) regularity compared to the case of noise diffused inside the domain. This is due to the large amplitude of the fluctuations of the solutions close to the boundary. In particular, in the case of Dirichlet boundary conditions the solution is only a distribution. This allowed us to treat only a restricted class of nonlinearities, exploiting specific properties of the heat semigroup and studying carefully the blow-up of the solution close to the boundary. For some results in this direction, the reader is referred to [2, 12, 23, 34]. On the contrary, in the last few years, maximal $L^{p}$ regularity techniques provided new ideas to treat partial differential equations with white noise Neumann boundary conditions with more severe nonlinearities such as those coming from fluid dynamical models. Indeed, some results on the global and local well-posedness of the 2D Navier-Stokes equations and the 3D primitive equations with boundary noise perturbations of Neumann time have been proven in [1] and [10], respectively. Besides the physical interests in studying the Navier-Stokes equations with boundary noise in virtue of its connection with the Couette flow (see also below for further motivations), the present manuscript also aims at (partially) filling the gap in the literature between Dirichlet and Neumann type boundary conditions for fluid dynamical models.
Throughout the manuscript, we fix a finite time horizon $T>0$ and we let $\mathcal{O}=\mathbb{T}\times(0,a)$ where $\mathbb{T}$ is the one-dimensional torus and $a>0$. Moreover, we denote by
(1.1) $\Gamma_{b}=\mathbb{T}\times\{0\}\quad\text{ and }\quad\Gamma_{u}=\mathbb{T}\times\{a\},$
the bottom and the upper part of the boundary of $\mathcal{O}$, respectively.
In this paper we are interested in the global well-posedness and the interior regularity of the 2D Navier-Stokes equations with boundary noise for the unknown velocity field $u(t,\omega,x,z)=(u_{1},u_{2}):(0,T)\times\Omega\times\mathcal{O}\to{\mathbb{R}}^{2}$ and pressure $P:(0,T)\times\Omega\times\mathcal{O}\to{\mathbb{R}}$, formally written as
@@
According to [33, 50, 51], see also the discussion in the introduction of [1], the geometry considered in (1.2) can be seen as an idealization of the ocean dynamics (more precisely, a vertical slice of the ocean). The model (1.2), describes a Couette flow, namely a viscous fluid in the space between two surfaces, one of which is moving tangentially relative to the other. The relative motion of the surfaces imposes a shear stress on the fluid and induces the flow. Let us recall that the onset of turbulence is often related to the randomness of background movement [45]. Moreover, according to [52, Chapter 3] in any turbulent flow there are unavoidably perturbations in boundary conditions and material properties. We model these features by the noise term $g\,\dot{W}^{\mathcal{H}}$. As introduced by Kolmogorov in [39], fractional Brownian motion can be thought of as a model for turbulence. Moreover, to describe turbulence in 3D fluids, models of random vortex filaments have been introduced in [27]. These have been analyzed for fractional Brownian motion with $\mathcal{H}>1/2$ in [49] and $\mathcal{H}<1/2$ in [28].
@@ 1.1. Main result
We begin by introducing some notation. Throughout this manuscript, we consider a complete filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbf{P})$, a separable Hilbert space $U$. We say that a process $\Phi$ is $\mathcal{F}$-progressive measurable if $\Phi|_{(0,t)\times\Omega}$ is $\mathcal{F}_{t}\times\mathcal{B}((0,t))$-measurable for all $t>0$, where $\mathcal{B}$ denotes the Borel $\sigma$-algebra. We postpone to subsection 1.3 the relevant notation on function spaces. On the noise $W^{\mathcal{H}}$ we enforce the following
@@
Note that 1.1 is consistent with the results obtained in [22] for the stochastic heat equation with Dirichlet fractional noise. The reader is referred to 1.4 for the case of a time-dependent $g$. Following [21, Chapter 15] and [19], we look for solutions to (1.2) in the form
$u=w_{g}+v,$
@@
As discussed in [21, Chapter 13], if $g$, $u_{\mathrm{in}}$, $W^{\mathcal{H}}(t)$ were regular enough, then $u=v+w_{g}$ would be a classical solution of the Navier-Stokes equations with non-homogeneous boundary conditions (1.2).
Next, we introduce the class of solutions we are going to consider. To motivate them, let us first discuss the regularity of $w_{g}$. It is well-known that, in the case of Dirichlet-type boundary conditions, the solution of a linear problem with boundary noise and $\mathcal{H}=\frac{1}{2}$ is a distribution which blows-up close to the boundary, see [2, 20], and the same holds also in case of $\mathcal{H}\neq\frac{1}{2}$, see [12]. Therefore, we cannot expect that the mild solution of (1.3) has arbitrarily good integrability properties as in [1, 10]. This has drastic consequences in our analysis. As we will show in 3.1, we have, $\mathbf{P}-a.s.$,
@@
Let us stress that $q_{\mathcal{H}}<2$ and $\lim_{\mathcal{H}\downarrow 3/4}q_{\mathcal{H}}\leq 1$ even if $s=0$. As we will see below, this fact creates major difficulties in our analysis of the auxiliary Navier-Stokes equations (1.4). In particular, $w_{g}\otimes w_{g}\in C([0,\infty);L^{q}(\mathcal{O};{\mathbb{R}}^{2}))$ $\mathbf{P}-a.s.$ and, from parabolic regularity, the best regularity we can hope for is $v\in L^{p}([0,\infty);H^{1,q}(\mathcal{O};{\mathbb{R}}^{2}))$ $\mathbf{P}-a.s.$ for all $p<\infty$. Thus, in general,
@@
Therefore, $v$ is a solution of the Navier-Stokes equations with infinite energy and the argument used in [1] does not work. The case of infinite energy solutions of 2D Navier-Stokes equations already appeared in the literature [11, 31]. In [31] the unboundedness of the energy is due to a rough initial data $u_{0}\not\in L^{2}$ while in [11] to a rough forcing term $f\not\in L^{2}(0,T;H^{-1})$ acting on the bulk. Our case does not fit in any of the above situations due to the presence of transport-type terms depending on the $w_{g}$ in (1.4) and the fact that we are working on domains. For this reason, our proofs rely on different methods. For details, the reader is referred to the text before 1.4.
In light of the previous discussion, we are now ready to define solutions to (1.2). Below, we set $A:B=\sum_{i,j=1}^{2}A^{i,j}B^{i,j}$ for two matrices $A$ and $B$ and $\mathbb{L}^{q}$ the image of $L^{q}(\mathcal{O};{\mathbb{R}}^{2})$ via the Leray projection $\mathbb{P}$ defined rigorously in subsection 2.1.
@@
  • ($q$-solution) A progressively measurable process $u$ with $\mathbf{P}-a.s.$ paths in $L^{2q^{\prime}}(0,T;\mathbb{L}^{2q})$, is a pathwise weak $q$-solution of (1.2) if $u=v+w_{g}$ and $v$ satisfies for all divergence-free $\varphi\in C^{\infty}(\mathcal{O};{\mathbb{R}}^{2})$ such that $\varphi=0$ on $\Gamma_{b}\cup\Gamma_{u}$ and a.e. $t\in(0,T)$,
    $\int_{\mathcal{O}}v(x,t)\varphi(x)\,\mathrm{d}x-\int_{\mathcal{O}}u_{\mathrm{in}}(x)\varphi(x)\,\mathrm{d}x$
  • (unique $q$-solution) A $q$-solution $u$ to (1.2) is said to be a unique if for any other $q$-solution $\widetilde{u}$ we have $u=\widetilde{u}$ a.e. on $[0,T]\times\Omega$.
@@
Before stating our main result, let us first comment on the above definition. Due to the argument below (1.4), one cannot expect solutions to (1.2) with integrability in space larger or equal to $2q_{\mathcal{H}}$. Furthermore, unique solution of (1.2) are independent of the choice of $q\in(1,q_{\mathcal{H}})$. Such independence is expected from solutions to (1.2) in light of (1.5). Finally, let us discuss the regularity class chosen to define $q$-solutions. Since $\mathcal{O}$ is two-dimensional, the space $L^{2q^{\prime}}(0,T;L^{2q}(\mathcal{O};{\mathbb{R}}^{2}))$ has Sobolev index given by (keeping in mind the parabolic scaling)
@@
    There exists a unique solution of (1.2) in the sense of 1.2 with paths in
@@
The proof of Theorem 1.3(1) and (2) are given subsection 3.3 and subsection 4.2, respectively. Routine extensions of the above are commented in 1.4 below.
@@
where $N$ depends only on $q$. The terms $\{v_{i}\}_{i\in\{0,\dots,N-1\}}$ are defined inductively solving homogeneous Stokes equations with forcing having mixed regularity in space and time, such that the regularity in space increases in $i$ while the regularity in time decreases in $i$. On the contrary $\overline{v}$ is a Leray-type solution of the remainder equation. In particular, $N$ is chosen large enough such that the equation for $\overline{v}$ has a forcing in $L^{2}(0,T;H^{-1})$ and therefore is regular enough to prove the existence and uniqueness of Leray solutions, see Theorem 2.5 below. The reader is referred to subsection 3.2 for further discussions.
The interior regularity of $u$ in Theorem 1.3(1) is treated considering again the splitting $u=v+w_{g}$ introduced above. The interior regularity of $w_{g}$ can be proved similarly to the linear part of [1]. On the contrary, the low regularity of $v$ does not allow us to study directly its interior regularity by Serrin’s argument as in [1]. For this reason, we rely on the splitting (1.7) analysed in subsection 3.2 to study the well-posedness of (1.4). Combining maximal $L^{p}$ regularity techniques for studying the interior regularity of the $v_{i}$’s, an induction argument and a Serrin argument for treating the interior regularity of $\overline{v}$, we obtain the required regularity of $v$. According to [58] (see also [41, Section 13.1]), it seems not possible to gain higher-order interior time-regularity for the Navier-Stokes problem. This fact is in contrast to the case of the heat equation with white noise boundary conditions, see [12]. The reason behind this is the presence of the unknown pressure $P$ which, due to its non-local nature, provides a connection between the interior and the boundary regularity where the noise acts.
To conclude, let us note that, in contrast to [11, 31], we employ a different splitting scheme to prove existence due to the presence of the transport-type terms originated by $w_{g}$. Moreover, the number of splitting $N$ depends on how much the Sobolev index of the space $C([0,T];L^{2q}(\mathcal{O};{\mathbb{R}}^{2}))$, i.e. $-\frac{1}{q}$, is far from the critical threshold $-1$. In particular, $N\to\infty$ as $q\downarrow 1$. As commented above, such a splitting is also convenient when proving the interior regularity for $u$ which was not addressed in the above-mentioned works.
@@
  • (The case of bounded domains) If $\mathcal{O}$ is replaced by a smooth $C^{2}$-bounded domain in ${\mathbb{R}}^{2}$. However, we prefer to keep the same geometry of [1] for two reasons. Firstly, and more importantly, as discussed in section 1, the model considered has a clear physical interpretation. Secondly, in this way, we can easily compare our results, techniques and assumptions with those of [1].
@@
  • (Time-dependent $g$) The term $g$ in the boundary noise $g\dot{W}^{\mathcal{H}}$ depends on time as long as it is progressively measurable and the corresponding process $w_{g}$ satisfies (1.5) holds.

### v2

@@ Acknowledgements.
The authors thank Robert Denk and Tim Seitz for several useful discussions. Finally, the authors thank the anonymous referees for helpful comments which improved the paper from its initial version.
@@ 1. Introduction
In many situations occurring in applied sciences, noise can affect the evolution of a system only through the boundary of a region where the system evolves. Such phenomena can be modeled via partial differential equations with boundary noise, as introduced by Da Prato and Zabczyck in the seminal paper [21]. Such a description presents several issues from a mathematical viewpoint. Indeed, nowadays it is well-known that in the one dimensional case, the solution of the heat equation with white noise Dirichlet or Neumann boundary conditions has low (space) regularity compared to the case of noise diffused inside the domain. This is due to the large amplitude of the fluctuations of the solutions close to the boundary. In particular, in the case of Dirichlet boundary conditions the solution is only a distribution. This allowed to treat only a restricted class of nonlinearities, exploiting specific properties of the heat semigroup and studying carefully the blow-up of the solution close to the boundary. For some results in this direction, the reader is referred to [2, 12, 25, 36]. On the contrary, partial differential equations with white noise Neumann boundary conditions with more severe nonlinearities have been considered in [14, 61], and in the last few years, maximal $L^{p}$ regularity techniques provided new ideas to treat some of those arising in fluid dynamics. Indeed, some results on the global and local well-posedness of the 2D Navier-Stokes equations and the 3D primitive equations with boundary noise perturbations of Neumann type have been proven in [1] and [10], respectively. Besides the physical interests in studying the Navier-Stokes equations with boundary noise in virtue of its connection with the Couette flow (see also below for further motivations), the present manuscript also aims at (partially) filling the gap in the literature between Dirichlet and Neumann type boundary conditions for fluid dynamical models.
Throughout the manuscript, we fix a finite time horizon $T>0$ and consider a spatial domain $\mathcal{O}=\mathbb{T}\times(0,a)$ where $\mathbb{T}$ is the one-dimensional torus and $a>0$. We further define the lower and upper parts of the boundary of $\mathcal{O}$ by
(1.1) $\Gamma_{b}=\mathbb{T}\times\{0\}\quad\text{ and }\quad\Gamma_{u}=\mathbb{T}\times\{a\}.$
In this work, we focus on the global well-posedness and interior regularity of the two-dimensional Navier–Stokes equations with fractional boundary noise. The unknowns are the velocity field $u(t,\omega,x,z)=(u_{1},u_{2}):(0,T)\times\Omega\times\mathcal{O}\to{\mathbb{R}}^{2}$ and the pressure $P:(0,T)\times\Omega\times\mathcal{O}\to{\mathbb{R}}$, which formally satisfy the system
@@
According to [35, 52, 53], see also the discussion in the introduction of [1], the geometry considered in (1.2) can be seen as an idealization of the ocean dynamics (more precisely, a vertical slice of the ocean). The model (1.2), describes a Couette flow, namely a viscous fluid in the space between two surfaces, one of which is moving tangentially relative to the other (see also Remark 1.4). The relative motion of the surfaces imposes a shear stress on the fluid and induces the flow. Let us recall that the onset of turbulence is often related to the randomness of background movement [47]. Moreover, according to [54, Chapter 3] in any turbulent flow there are unavoidably perturbations in boundary conditions and material properties. We model these features by the noise term $g\,\dot{W}^{\mathcal{H}}$. As introduced by Kolmogorov in [41], fractional Brownian motion can be thought of as a model for turbulence. Moreover, to describe turbulence in 3D fluids, models of random vortex filaments have been introduced in [29]. These have been analyzed for fractional Brownian motion with $\mathcal{H}>1/2$ in [51] and $\mathcal{H}<1/2$ in [30].
@@ 1.1. Main result
We begin by introducing some basic notation. Throughout this manuscript, we work on a complete filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},\mathbf{P})$ and consider a separable Hilbert space $U$. A process $\Phi$ is said to be $\mathcal{F}$-progressive measurable if, for every $t>0$, the restriction $\Phi|_{(0,t)\times\Omega}$ is measurable with respect to $\mathcal{F}_{t}\otimes\mathcal{B}((0,t))$, where $\mathcal{B}$ denotes the Borel $\sigma$-algebra. Further notation concerning function spaces is deferred to Section 1.3. On the noise $W^{\mathcal{H}}$ we enforce the following
@@
Note that 1.1 is consistent with the results obtained in [23] for the stochastic heat equation with Dirichlet fractional noise. The reader is referred to Remark 1.4 for the case of a time-dependent $g$. Following [22, Chapter 15] and [20], we construct solutions to (1.2) by the splitting
(1.3) $u=w_{g}+v,$
@@
As discussed in [22, Chapter 13], if $g$, $u_{\mathrm{in}}$, $W^{\mathcal{H}}(t)$ were sufficiently regular, then $u=v+w_{g}$ would be a classical solution of the Navier-Stokes equations with non-homogeneous boundary conditions (1.2).
Next, we introduce the class of solutions we are going to consider. To motivate them, let us first discuss the regularity of $w_{g}$. It is well-known that, in the case of Dirichlet-type boundary conditions, the solution of a linear problem with boundary noise and $\mathcal{H}=\frac{1}{2}$ is a distribution which blows-up close to the boundary, see [2, 21], and the same holds also in case of $\mathcal{H}\neq\frac{1}{2}$, see [12]. Therefore, we cannot expect that the mild solution of (1.4) has arbitrarily good integrability properties as in [1, 10]. This has drastic consequences in our analysis. As we will show in Proposition 3.1, we have, $\mathbf{P}-a.s.$,
@@
Let us stress that $q_{\mathcal{H}}<2$ and $\lim_{\mathcal{H}\downarrow 3/4}q_{\mathcal{H}}{=}1$ even if $s=0$. As we will see below, this fact creates major difficulties in our analysis of the auxiliary Navier-Stokes equations (1.5). In particular, $w_{g}\otimes w_{g}\in C([0,\infty);L^{q}(\mathcal{O};{\mathbb{R}}^{2}))$ $\mathbf{P}-a.s.$ and, from parabolic regularity, the best regularity we can hope for is $v\in L^{p}([0,\infty);H^{1,q}(\mathcal{O};{\mathbb{R}}^{2}))$ $\mathbf{P}-a.s.$ for all $p<\infty$. Thus, in general,
@@
Therefore, $v$ is a solution of the Navier-Stokes equations with infinite energy and the argument used in [1] does not work. The case of infinite energy solutions of 2D Navier-Stokes equations already appeared in the literature [11, 33]. In [33] the unboundedness of the energy is due to a rough initial data $u_{0}\not\in L^{2}$ while in [11] to a rough forcing term $f\not\in L^{2}(0,T;H^{-1})$ acting on the bulk. Our case does not fit in any of the above situations due to the presence of transport-type terms depending on the $w_{g}$ in (1.5) and the fact that we are working on domains. For this reason, our proofs rely on different methods. For details, the reader is referred to the text before Remark 1.4.
In light of the previous discussion, we are now ready to define solutions to (1.2). Below, we set $A:B=\sum_{i,j=1}^{2}A^{i,j}B^{i,j}$ for two matrices $A$ and $B$ and $\mathbb{L}^{q}$ the image of $L^{q}(\mathcal{O};{\mathbb{R}}^{2})$ via the Leray projection $\mathbb{P}$ defined rigorously in Section 2.1.
@@
  • ($q$-solution) A progressively measurable process $u$ with $\mathbf{P}-a.s.$ paths in $L^{2q^{\prime}}(0,T;\mathbb{L}^{2q})$, is a pathwise weak $q$-solution of (1.2) if for all divergence-free $\varphi=(\varphi_{1},\varphi_{2})\in C^{\infty}(\mathcal{O};{\mathbb{R}}^{2})$ such that $\varphi=0$ on $\Gamma_{b}\cup\Gamma_{u}$ and a.e. $t\in(0,T)$,
    $\int_{\mathcal{O}}u(x,t)\varphi(x)\,\mathrm{d}x-\int_{\mathcal{O}}u_{\mathrm{in}}(x)\varphi(x)\,\mathrm{d}x$
    $\qquad\qquad=\int_{0}^{t}\int_{\mathcal{O}}\big{(}u\cdot\Delta\varphi+[u\otimes u]:\nabla\varphi\big{)}\,\mathrm{d}x\,\mathrm{d}r-{\langle g,\partial_{2}\varphi_{1}\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}W^{\mathcal{H}}_{t}}.$
  • (unique $q$-solution) A $q$-solution $u$ to (1.2) is said to be a unique solution if for any other $q$-solution $\widetilde{u}$ we have $u=\widetilde{u}$ a.e. on $[0,T]\times\Omega$.
@@
Before stating our main result, let us first comment on the above definition. Due to the argument below (1.5), one cannot expect solutions to (1.2) with integrability in space larger or equal to $2q_{\mathcal{H}}$. Furthermore, even if solutions are constructed by (1.3), the uniqueness class is independent of this splitting. Moreover, the unique solution of (1.2) is independent of the choice of $q\in(1,q_{\mathcal{H}})$. Such independence is expected from solutions to (1.2) in light of (1.6). Finally, let us discuss the regularity class chosen to define $q$-solutions. Since $\mathcal{O}$ is two-dimensional, the space $L^{2q^{\prime}}(0,T;L^{2q}(\mathcal{O};{\mathbb{R}}^{2}))$ has Sobolev index given by (keeping in mind the parabolic scaling)
@@
    There exists a unique solution of (1.2) in the sense of Definition 1.2 with paths in
@@
The proof of Theorem 1.3(1) and (2) are given in Section 3.3 and Section 4.2, respectively. Routine extensions of the above are commented in Remark 1.4 below.
@@
where $N$ depends only on $q$. The terms $\{v_{i}\}_{i\in\{0,\dots,N-1\}}$ are defined inductively solving homogeneous Stokes equations with forcing having mixed regularity in space and time, such that the regularity in space increases in $i$ while the regularity in time decreases in $i$. On the contrary $\overline{v}$ is a Leray-type solution of the remainder equation. In particular, $N$ is chosen large enough such that the equation for $\overline{v}$ has a forcing in $L^{2}(0,T;H^{-1})$ and therefore is regular enough to prove the existence and uniqueness of Leray solutions, see Theorem 2.6 below. The reader is referred to Section 3.2 for further discussions.
The interior regularity of $u$ in Theorem 1.3(2) is treated considering again the splitting $u=v+w_{g}$ introduced above. The interior regularity of $w_{g}$ can be proved similarly to the linear part of [1]. On the contrary, the low regularity of $v$ does not allow us to study directly its interior regularity by Serrin’s argument as in [1]. For this reason, we rely on the splitting (1.8) analysed in Section 3.2 to study the well-posedness of (1.5). Combining maximal $L^{p}$ regularity techniques for studying the interior regularity of the $v_{i}$’s, an induction argument and a Serrin argument for treating the interior regularity of $\overline{v}$, we obtain the required regularity of $v$. As shown in [60] (see also [43, Section 13.1]) and similarly to [1], higher-order interior time regularity does not appear to be attainable for our Navier–Stokes problem with stochastic boundary conditions. This stands in contrast to the case of the heat equation with white noise boundary conditions, as studied in [12]. The underlying reason for this phenomenon is the presence of the unknown pressure $P$, whose non-local nature creates a link between interior and boundary regularity, where the noise acts.
To conclude, let us point out that, in contrast to [11, 33], we employ a different splitting scheme to prove existence due to the presence of the transport-type terms originated by $w_{g}$. Moreover, the number of splitting $N$ depends on how much the Sobolev index of the space $C([0,T];L^{2q}(\mathcal{O};{\mathbb{R}}^{2}))$, i.e. $-\frac{1}{q}$, is far from the critical threshold $-1$. In particular, $N\to\infty$ as $q\downarrow 1$. As commented above, such a splitting is also convenient when proving the interior regularity for $u$ which was not addressed in the above-mentioned works.
@@
  • (Bounded domains) If $\mathcal{O}$ is replaced by a smooth $C^{2}$-bounded domain in ${\mathbb{R}}^{2}$. However, we prefer to keep the same geometry of [1] for two reasons. Firstly, and more importantly, as discussed in Section 1, the model considered has a clear physical interpretation. Secondly, in this way, we can easily compare our results, techniques and assumptions with those of [1].
@@
  • (Time-dependent $g$) The term $g$ in the boundary noise $g\dot{W}^{\mathcal{H}}$ depends on time as long as it is progressively measurable and the corresponding process $w_{g}$ satisfies (1.6).
  • (Full stochastic Couette flow) More general boundary conditions like
    $\begin{cases}u_{1}=U_{up}+g\dot{W}^{\mathcal{H}}\quad&\text{on }(0,T)\times\Gamma_{u},\\ u_{1}=U_{b}\quad&\text{on }(0,T)\times\Gamma_{b}.\end{cases}$
    can be considered for sufficiently smooth velocity fields $U_{up},U_{b}$ such that the corresponding process $w_{g}$ satisfies (1.6). For example the case $U_{up},U_{b}\in L^{2}((0,T)\times\mathbb{T})$ can be treated. The above can be seen as a Couette flow with uncertainty on the velocity of one of the two surfaces.